OpenStudy (anonymous):

Simplify. Remember to use absolute-value notation when necessary. If a root cannot be simplified, state this.

OpenStudy (anonymous):

\[\sqrt[12]{-10}^{12}\]

OpenStudy (anonymous):

this is what I got but i was wrong 12((-10) □(1/2))12 12(-10)6 (12*(-10) ┤ (-10)(-10)(-10)(-10)(-10)) Multiply to simplify (12*1000000) Multiply by 12 by 100000 then you will get (12000000) Answer is 12000000

OpenStudy (anonymous):

Is your initial expression\[\sqrt[12]{(-10)^{12}}\]?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Then your answer is |10| which can also be written +-10 because you have an even-powered root. So, a negative squared or "twelved" will be positive. And the 10 in the answer comes from the exponent being 12/12 = 1

OpenStudy (anonymous):

How would you set that up, because that is what I was having a problem with

OpenStudy (anonymous):

np\[\sqrt[12]{(-10)^{12}} = \sqrt{(-10)^{2}} = \sqrt{100} = \pm10\]

OpenStudy (anonymous):

WOW i knew I was wrong but not that wrong

OpenStudy (anonymous):

Thank you for your help! And if I set it up like this i will be correct

OpenStudy (anonymous):

These things can be tricky!

OpenStudy (anonymous):

Yes. Good luck to you and thx for the recognition!

OpenStudy (anonymous):

thank you for your help